OpenStudy (anonymous):

Calculus differential equations: If dx/dt= (2t+1)/(2x) and x(1)=3 find x(0) Please help!

OpenStudy (anonymous):

I'm not sure how I should start, should I try to separate and integrate?

OpenStudy (phi):

rewrite it as 2x dx = (2t+1) dt

OpenStudy (anonymous):

ok and then integrate each side... let me do that pretty basic so far, I'm just not sure what to do after that ah, give me a few seconds to integrate

OpenStudy (anonymous):

x^2= t^2 +t +C ?

OpenStudy (phi):

yes. now you need to find C use t=1 , x=3

OpenStudy (mathmale):

Looks good! Now use the initial condition, x(1)=3, to determine your constant of integration. In other words, take your own equation (above), and let t=1 and x=3. C will be the only variable left. You can calculate C from this equation.

OpenStudy (anonymous):

ok so sub in and C= 7

OpenStudy (phi):

now solve for x when t=0

OpenStudy (anonymous):

with this equation? x^2= t^2 +t +C x^2 = 0 +0 +7 x= sqrt7?

OpenStudy (phi):

yes

OpenStudy (anonymous):

wow thank you, you made it so easy! Would you help me with one more of this type please?